Circular No. 9108

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2010 A5 (LINEAR) An apparently asteroidal object reported by the LINEAR project (dis-

covery observation tabulated below) and placed on the Minor Planet Center’s ‘NEOCP’ webpage has been reported as cometary by other CCD astrometrists. R. Ligustri (Udine, Italy, remotely using a 0.32-m f/9 reflector at RAS Observatory, Moorook, Australia; Jan. 15.7 UT) writes that 24 stacked 60-s image show a coma of size ∼ 20′′. L. Buzzi (Varese, Italy, 0.60-m reflector; Jan. 16.2) reports that stacked images totalling 15 min of exposure time obtained in good seeing show a compact coma elongated in p.a. 300^{o}, with a tail visible in that direction for at least 20′′. K. Sárneczky (Konkoly Observatory, 1.02-m Ritchey-Chretien telescope; Jan. 16.2) notes that a 120-s co-added image shows a 25′′-long fan-shaped tail in p.a. 290^{o}. E. Guido and G. Sostero (remotely using a 0.25-m reflector near Mayhill, NM, U.S.A.; Jan. 16.4) write that fifteen co-added 120-s exposures show a diffuse coma nearly 12′′ in diameter, elongated ∼ 20′′ in p.a. 300^{o}. W. H. Ryan (Magdalena Ridge Observatory, 2.4-m reflector; Jan. 16.5) reports that R-band exposures taken by E. V. Ryan and himself show a coma with a tail in p.a. ∼ 280^{o}. H. Sato (Ota-ku, Tokyo, Japan, remotely using a 0.25-m reflector at RAS Observatory, near Mayhill, NM; Jan. 16.5) finds a 10′′ coma elongated toward p.a. 240^{o}. S. Foglia (using a 0.18-m reflector at Tzec Maun Observatory, Cloudcroft, NM, U.S.A., with P. Concari, G. Galli, and M. Tombelli; Jan. 16.5) reports that eight stacked 120-s images show the object to be diffuse with mag ∼ 17.1.

2010 UT α_{2000} δ_{2000} Mag. Observer
Jan. 14.45298 13^{h}26^{m}02.^{s}76 −15^{o}55′53′′.3 17.7 LINEAR

Further astrometry, the following preliminary elliptical orbital elements, and an ephemeris are given on MPEC 2010-B02.

T = 2010 Apr. 20.601 TT  ω = 307.^{o}527
e = 0.67268  Ω = 278.071 }2000.0
q = 1.70451 AU  i = 5.789
a = 5.20739 AU n^{o} = 0.082942 P = 11.9 years

COMET P/2010 A3 (HILL) Elliptical orbital elements from MPEC 2010-A85 for epoch 2010 Mar.

25.0 TT: T = 2010 Apr. 3.6804 TT, e = 0.732116, q = 1.621774 AU, i = 15^{o}.0270, ω = 41^{o}.2825, Ω = 64.8307 (equinox 2000.0), P = 14.9 years.

2010 January 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green